# Automated forest inventory: analysis of high-density airborne LiDAR point clouds with 3D deep learning

### 22 Dec 2023 v1

@@ 2.2.2 Network architecture
The point cloud segmentation network consists of a shared 3D feature extraction backbone, followed by three parallel prediction heads. The first head is dedicated to semantic segmentation. The two others extract two complementary embeddings of each point, which serve as the basis for tree instance clustering.
@@
The instance segmentation branch assigns point-wise individual tree IDs by clustering the outputs of the two corresponding prediction branches. One of them estimates, for each point, a 3D offset vector from the center of its tree instance, and is supervised with loss consisting of (i) the cosine distance between the true and predicted offset vectors and (ii) the $L_{1}$ distance between their endpoints. The 3D centre offset, advocated by several previous studies (Jiang et al., 2020; Chen et al., 2021a; Vu et al., 2022; Zhong et al., 2022) is essentially a learned version of the generalised Hough transform: perfect predictions would mean that the offsets contract each instance to a single point. The other branch maps each point to 5-dimensional embedding space, such that points on the same tree instance form clusters in that space. It is supervised with a contrastive loss function that aims to minimise distances between points on the same tree, and to maximise distances between points on different trees (Wang et al., 2019a). Notably, the two additional dimensions of the embedding space make it possible to represent further object properties beyond 3D geometric offsets.
The two embeddings are separately clustered to tree candidates: offsets are explicitly added to the point coordinates and the resulting, shifted points are clustered with simple region growing, with a threshold for the maximum distance. Clusters in the 5D embedding space, where distances do not have a direct geometric interpretation, are found with the mean-shift method. The two (redundant) sets of tree instance candidates are merged and filtered with another small neural network, called ScoreNet (Jiang et al., 2020), that predicts how well each instance candidate matches a ground truth (GT) tree instance (i.e., ScoreNet (Jiang et al., 2020)). That ScoreNet is a small 3D U-Net followed by max-pooling and a fully connected layer, and is trained to regress the maximal expected intersection over union (IoU) between an instance candidate and any of the actual trees. The candidates are then ordered by their score and pruned with greedy non-maximum suppression to obtain the final set of individual tree instances. All non-tree points are assigned the instance label $-1$. See (Xiang et al., 2023a) for further details.
@@ 2.3 Automated retrieval of tree parameters and stand structure
After segmenting a 3D point cloud, important individual tree attributes and stand structure characteristics can be extracted with straightforward geometric computations. These biophysical attributes are the actual variables of interest for the forest inventory, because they typically serve as input for management decisions and ecological analysis. Compared to field measurements, tree parameters retrieved from point clouds allow for complete and consistent coverage of much larger areas. For some parameters (e.g., tree height), LiDAR retrievals are also the most accurate method; while for others (e.g., DBH) they are not as accurate as field measurements, and one has to accept a lower reliability of the individual measurement in exchange for the better coverage. In the present study, we extract the following exemplary forest attributes: at plot level, a digital terrain model (DTM) of the plot and its stand density; and for each individual tree the height, crown diameter, crown volume, crown volume of live branches, DBH and location.
@@ 2.4.1 Metrics for point cloud segmentation
Table 2: Metrics used to evaluate individual tree segmentation. In machine learning, including computer vision, several of these metrics are also common, but have different names. Metrics printed in bold font are the ones used in the present study.
@@
User’s accuracy (UAdetect, Detection rate)
Correctness
UAdetect=TP/(TP+FP) Precision (p)
Accuracy index (AI) AI = (TP-FP)/N \
Matching score (M) M = TP/(N+FP) \
@@
When comparing the literature, we found that forest researchers (Yin and Wang, 2016; Liang et al., 2018) and machine learning professionals, including those specialising in computer vision (Gu et al., 2022; Wang et al., 2019a), often use different names for the same quality metric. To make this explicit and avoid misinterpretations, we list them in Table 2. It shows that both fields differ in naming conventions while conveying similar evaluation criteria. In the Table 2, $N$ is the total number of reference trees, TP is the number of true positives (correctly detected trees), FN is number of reference trees that are not detected, FP is number of false positives (detections not correspondiong to any reference tree), and TP+FP thus equals the total number of trees predicted by the system.
A subtle source of potential differences is the procedure used to match predicted trees to reference trees. In forest research this is frequently done based on geometric distances, i.e., each reference tree is matched to the detection that closest to it in terms of (stem) location (Yan et al., 2020; Huo et al., 2022; Hao et al., 2022). Sometimes additional exclusion criteria are employed to ensure the detection is sufficiently close, has similar height, etc. A related variant is to consider a detection correct if its location falls within the crown boundary of the reference tree (Chen et al., 2022). Some studies on individual tree detection match by IoU score between projected 2D crown polygons (Dietenberger et al., 2023). In our study we follow the standard procedure in computer vision, which reflects the 3-dimensional nature of the problem. Detections are matched with reference trees at the point level, and matches are made based on the IoU score between the two point sets. Predictions that do not reach an IoU of at least 0.5 with any reference tree are counted as false positives.
@@ 2.5 Implementation details
All code for our study was written in Python. All experiments were conducted on a machine with 8-core Intel CPU, 8 GB of memory per core, and one Nvidia Titan RTX GPU with 24 GB of on-board memory. The manual feature extraction code is a slightly modified version of the code published by Guinard and Landrieu (2017). The point cloud segmentation network was constructed with the help of the Torch-Point3D library (Chaton et al., 2020). All hyper-parameters for HDBSCAN filtering were determined by grid search on the validation set.
@@ 3 Results
A series of experiments was conducted to examine whether various loss designs, weighting strategies and data augmentations could enhance segmentation performance for forest scenes. The quantitative results for individual tree segmentation and semantic segmentation, across different settings, are summarized in Table 3. The classification results for each semantic class, as well as the binary tree vs. non-tree classification accuracy, are shown in Table 4. The basic setting is the panoptic segmentation network we developed in earlier work (Xiang et al., 2023a), which we expanded to distinguish five semantic classes instead of only two (tree and non-tree). The other tested settings (Tables 3 and 4) are obtained by applying a single modification to the base setting.
Table 3: Instance segmentation and semantic segmentation results under different settings. Bold values in each column indicate the best results.
@@
Completeness Omission error Commission error F-score oAcc mAcc mIoU
Basic setting 79.3 20.8 21.2 79.0 92.6 81.2 73.0
+ binary semantic loss 83.0 17.0 21.6 80.6 92.5 80.4 72.4
+ class weights 78.9 21.1 22.5 78.2 90.9 83.9 70.7
+ height weights 80.8 19.2 21.4 79.7 91.8 78.4 70.4
+ region weights 80.5 19.5 20.0 80.2 92.3 79.0 71.4
+ intensity 78.9 21.1 14.7 82.0 93.0 78.9 72.7
+ return number 81.7 18.3 13.5 84.1 92.9 81.6 73.7
+ scan angle rank 81.7 18.3 18.5 81.6 93.2 81.8 74.3
+ manually-extracted features 81.4 18.6 18.1 81.7 93.6 82.6 75.7
+ elastic distortion and dropout 83.3 16.7 16.5 83.4 92.6 80.5 72.4
+ TreeMix 82.0 18.0 11.7 85.1 93.0 80.8 73.5
@@
+ manually-extracted features 89.1 78.7 55.6 94.3 60.7 75.7 94.2 97.3 95.7
+ elastic distortion and dropout 85.8 73.9 50.8 93.9 57.8 72.4 92.6 96.6 94.6
@@
The tested setting can be categorized into three groups. The first group modifies the loss function or the sampling of training examples, i.e., the available input data is not modified, but used in a different manner. One way to do this is to add explicit supervision for the binary tree vs. non-tree task. At first glance this may appear superfluous, since the binary reference is derived by simply aggregating the fine-grained labels (stems, live branches and dead branches into the tree class, ground and low vegetation into the non-tree class). Still the aggregation adds information that otherwise is unavailable to the network, namely the hierarchical structure of the label space. E.g., as soon as a point is part of a tree, it can no longer be confused with the low vegetation, hence the classifier can invest its capacity into difficult decisions within the tree class, such as separating live from dead branches. Indeed, we empirically find that adding the binary labels increases the IoU for tree vs. non-tree by 1.8 percent points (Table 4). A more correct set of tree points, in turn, better supports instance discrimination, so that the F-score for individual tree segmentation also improves by 1.6 percent points (see Table 3). The price to pay is a slightly higher confusion between the non-tree classes ground and low vegetation.
Also in this group fall different sampling and reweighting strategies. Sampling based on class frequency significantly improves the segmentation of the under-represented stem class (+3.4 pp), and also of the dead branches (+1.2 pp). However the associated accuracy loss for the ground (--2.3 pp) hurts overall performance. Reweighting based on point height is meant to compensate the sampling bias of the ALS recording geometry, vertically through the canopy. Reweighting by the regional class distribution is a coarser version of class balancing, based on regional forest composition. The latter two strategies, however, both result in a moderate decline in IoU, despite small gains in regions with little training data.
@@
The third group are tree-specific data augmentation strategies to synthetically increase the size and diversity of the training set. We have experimented with two approaches: elastic distortion with dropout, and TreeMix. The idea behind the use of elastic distortion is to better account for curved and non-vertical stems, which are too rare to be learned well, but may still occur in the test data. The reasoning behind dropout, i.e., removing points from the training examples, is to simulate points lost due to occlusion, scattering, etc., and make the network more robust against missing data. These low-level augmentation methods increase the instance segmentation F-score by 4.4 pp, but slightly reduce the semantic segmentation performance. TreeMix is a more informed augmentation strategy, where samples are synthesised by mixing semantically meaningful entities (i.e., trees) from different samples. It achieves the best single tree segmentation result of 85.1% F-score (+6.1 pp), while also improving semantic segmentation.
@@
Figure 3: Confusion matrix between the five semantic categories, across all points of the test set.
@@
In addition, NIBIO, CULS, and SCION are pure coniferous or conifer-dominated forests. Coniferous trees are generally straighter and have slimmer crowns with smaller branches. On the contrary, RMIT is native dry sclerophyll Eucalypt forest, where eucalyptus trees have few, but thick branches and expansive crowns. TUWIEN is a deciduous-dominated alluvial forest, where trees have many branches and low crowns, varying greatly in size. Expectedly, our results show that segmenting tall and straight, slender trees is easier; even when they are located close to each other (Table 5 and Figure 4). Whereas for trees with dispersed crowns and complex branch structure, individual tree segmentation is more challenging. A further reason for this discrepancy the limited diversity of training data for the forest characteristics of RMIT and TUWIEN, which each only have a single plot of training data.
Table 5: Individual tree segmentation results for each plot in the test set.
@@
(Commission rate)
F-score
1 (CULS) 20 20 23 00000020 (100.0%) 000000 (0.0%) 0000003 (13.0%) 93.0%
2 (NIBIO) 37 26 26 26 (70.3%) 11 (29.7%) 0 (0.0%) 82.5%
3 (NIBIO) 30 23 29 29 (96.7%) 1 (3.3%) 0 (0.0%) 98.3%
4 (NIBIO) 27 21 26 26 (96.3%) 1 (3.7%) 0 (0.0%) 98.1%
5 (NIBIO) 20 15 18 18 (90.0%) 2 (10.0%) 0 (0.0%) 94.7%
6 (NIBIO) 28 18 29 26 (92.9%) 2 (7.1%) 3 (10.3%) 91.2%
7 (NIBIO) 19 19 18 16 (84.2%) 3 (15.8%) 2 (11.1%) 86.5%
8 (RMIT) 64 27 54 41 (64.1%) 23 (35.9%) 13 (24.1%) 69.5%
9 (SCION) 25 25 25 23 (92.0%) 2 (8.0%) 2 (8.0%) 92.0%
10 (SCION) 18 18 15 15 (83.3%) 3 (16.7%) 0 (0.0%) 90.9%
11 (TUWIEN) 35 35 37 25 (71.4%) 10 (28.6%) 12 (32.4%) 69.4%

### 23 Feb 2024 v2

@@ 2.2.2 Network architecture
The point cloud segmentation network consists of a shared 3D feature extraction backbone, followed by three parallel prediction heads Δ=Δ Δ=Δ Δ=Δ Δ=Δ (Figure 2). The first head is dedicated to semantic segmentation. Δ=Δ Δ=Δ Δ=Δ Δ=Δ To address the complexity of segmenting trees in densely packed forests, it proved beneficial to use two heads to extract two complementary embeddings of each point, which serve as the basis for tree instance clustering.
@@
The instance segmentation branch assigns point-wise individual tree IDs by clustering the outputs of the two corresponding prediction branches. One of them estimates, for each point, a 3D offset vector from the center of its tree instance, and is supervised with loss consisting of (i) the cosine distance between the true and predicted offset vectors Δ=Δ Δ=Δ Δ=Δ Δ=Δ , to align the directions of the predicted and true offset vectors, irrespective of object size and distance from the object center; and (ii) the $L_{1}$ distance between their endpoints Δ=Δ Δ=Δ Δ=Δ Δ=Δ to directly minimize the deviation from the true offset. The 3D centre offset Δ=Δ Δ=Δ Δ=Δ Δ=Δ has been advocated by several previous studies (Jiang et al., 2020; Chen et al., 2021a; Vu et al., 2022; Zhong et al., 2022) Δ=Δ Δ=Δ Δ=Δ Δ=Δ . Δ=Δ Δ=Δ Δ=Δ Δ=Δ Δ=Δ Δ=Δ Δ=Δ Δ=Δ A perfect prediction would mean that the offsets contract each instance to a single point. The other branch maps each point to a 5-dimensional embedding space, such that points on the same tree instance form clusters in that space. Δ=Δ Δ=Δ Δ=Δ Δ=Δ Using 5 dimensions is in line with existing literature (Wang et al., 2019a; Engelmann et al., 2020; He et al., 2020) and corroborated by our own experiments, where higher values did not produce significant accuracy gains. It is supervised with a contrastive loss function that aims to minimise distances between points on the same tree, and to maximise distances between points on different trees (Wang et al., 2019a). Notably, the two additional dimensions of the embedding space make it possible to represent further object properties beyond 3D geometric offsets. Δ=Δ Δ=Δ Δ=Δ Δ=Δ Empirically, using both the embedding and the direct offset prediction allows for to more accurate tree instance segmentation.
The two embeddings are separately clustered to tree candidates Δ=Δ Δ=Δ Δ=Δ Δ=Δ in unsupervised fashion: offsets are explicitly added to the point coordinates and the resulting, shifted points are clustered with simple region growing Δ=Δ Δ=Δ Δ=Δ Δ=Δ as recommended by (Zhao et al., 2021), with the threshold for the maximum distance Δ=Δ Δ=Δ Δ=Δ Δ=Δ set at 0.3 meters. Clusters in the 5D embedding space, where distances do not have a direct geometric interpretation, are found with the mean-shift method Δ=Δ Δ=Δ Δ=Δ Δ=Δ , as in several studies (Wang et al., 2019a; Lahoud et al., 2019). Mean-shift (Comaniciu and Meer, 2002) has only a single bandwidth parameter, which we always keep at an empirically determined value as 0.6, a setting we found to robustly work across city as well as forest scenes (Xiang et al., 2023b, a). In line with previous work, we find that conventional clustering as a post-process works better than learning cluster assignments within the neural network.
The two (redundant) sets of tree instance candidates are merged and filtered with another small neural network, called ScoreNet (Jiang et al., 2020) Δ=Δ Δ=Δ Δ=Δ Δ=Δ , that predicts how well each instance candidate matches a ground truth (GT) tree Δ=Δ Δ=Δ Δ=Δ Δ=Δ instance (i.e., ScoreNet (Jiang et al., 2020)). That ScoreNet is a small 3D U-Net followed by max-pooling and a fully connected layer, and is trained to regress the maximal expected intersection over union (IoU) between an instance candidate and any of the actual trees. The candidates are then ordered by their score and pruned with greedy non-maximum suppression to obtain the final set of individual tree instances. All non-tree points are assigned the instance label $-1$. See (Xiang et al., 2023a) for further details.
@@ 2.3 Automated retrieval of tree parameters and stand structure
After segmenting a 3D point cloud, important individual tree attributes and stand structure characteristics can be extracted with straightforward geometric computations Δ=Δ Δ=Δ Δ=Δ Δ=Δ (see Figure 1). These biophysical attributes are the actual variables of interest for the forest inventory, because they typically serve as input for management decisions and ecological analysis. Compared to field measurements, tree parameters retrieved from point clouds allow for complete and consistent coverage of much larger areas. For some parameters (e.g., tree height), LiDAR retrievals are also the most accurate method; while for others (e.g., DBH) they are not as accurate as field measurements, and one has to accept a lower reliability of the individual measurement in exchange for the better coverage. In the present study, we extract the following exemplary forest attributes: at plot level, a digital terrain model (DTM) of the plot and its stand density; and for each individual tree the height, crown diameter, crown volume, crown volume of live branches, DBH and location.
@@ 2.4 Evaluation metrics
Δ=Δ Δ=Δ Δ=Δ Δ=Δ For better readability, all metrics, including confusion matrices, mIoU scores and F-score, are shown as percentages, rounded to one decimal place.
@@ 2.4.1 Metrics for point cloud segmentation
Table 3: Δ=Δ Δ=Δ Δ=Δ Δ=Δ Metrics used to evaluate individual tree segmentation. In machine learning, including computer vision, several of these metrics are also in common use but have different names.
@@
When comparing the literature, we found that forest researchers (Yin and Wang, 2016; Liang et al., 2018) and machine learning professionals, including those specialising in computer vision (Gu et al., 2022; Wang et al., 2019a), often use different names for the same quality metric. To make this explicit and avoid misinterpretations, we list them in Table 3. It shows that both fields differ in naming conventions while conveying similar evaluation criteria. In the Table 3, $N$ is the total number of reference trees, TP is the number of true positives (correctly detected trees), FN is number of reference trees that are not detected, FP is number of false positives (detections not corresponding to any reference tree), and TP+FP thus equals the total number of trees predicted by the system.
A subtle source of potential differences is the procedure used to match predicted trees to reference trees. In forest research this is frequently done based on geometric distances, i.e., each reference tree is matched to the detection that closest to it in terms of (stem) location (Yan et al., 2020; Huo et al., 2022; Hao et al., 2022). Sometimes additional exclusion criteria are employed to ensure the detection is sufficiently close, has similar height, etc. A related variant is to consider a detection correct if its location falls within the crown boundary of the reference tree (Chen et al., 2022). Some studies on individual tree detection match by IoU score between projected 2D crown polygons (Dietenberger et al., 2023). In our study we follow the standard procedure in computer vision, which reflects the 3-dimensional nature of the problem. Δ=Δ Δ=Δ Δ=Δ Δ=Δ Our method matches predicted tree instances to reference trees based on the IoU score between their respective point sets. This approach ensures robustness against variations in tree and stem densities across different plots. Predictions that do not reach an IoU of at least 0.5 with any reference tree are counted as false positives.
To go beyond verifying the presence of tree instances and evaluate how precisely they delineated, we introduce the coverage metric, which quantifies the agreement with the ground truth instance boundaries. Denoting the set of ground truth trees $\{I_{i}^{\text{gt}},i\in\{1,...,N_{\text{gt}}\}\}$ and the set of predicted trees as $\{I_{j}^{\text{pre}},j\in\{1,...,N_{\text{pre}}\}\}$, we compare the two sets on a per-instance basis. For each ground truth tree we find the predicted instance with the highest IoU score, formally:
$\text{maxIoU}(I_{i}^{\text{gt}})=\max_{j=1}^{N_{\text{pre}}}\big{(}\text{IoU}(I_{i}^{\text{gt}},I_{j}^{\text{pre}})\big{)}\;.$ (1)
The coverage is then defined as the average over all reference trees:
$\text{Cov}=\frac{1}{N_{\text{gt}}}\sum_{i=1}^{N_{\text{gt}}}\text{maxIoU}(I_{i}^{\text{gt}})\;.$ (2)
@@ 2.5 Implementation details
All code for our study was written in Python Δ=Δ Δ=Δ Δ=Δ Δ=Δ , and is publicly available at https://github.com/bxiang233/ForAINet. All experiments were conducted on a machine with 8-core Intel CPU, 8 GB of memory per core, and one Nvidia Titan RTX GPU with 24 GB of on-board memory. The Δ=Δ Δ=Δ Δ=Δ Δ=Δ code for extracting hand-crafted geometric features is a slightly modified version of the code published by Guinard and Landrieu (2017). The point cloud segmentation network was constructed with the help of the Torch-Point3D library (Chaton et al., 2020). All hyper-parameters for HDBSCAN filtering were determined by grid search on the validation set.
@@ 3 Results
A series of experiments was conducted to examine whether various loss designs, weighting strategies and data augmentations could enhance segmentation performance for forest scenes. The quantitative results for individual tree segmentation and semantic segmentation, across different settings, are summarized in Table 4. The classification results for each semantic class, as well as the binary tree vs. non-tree classification accuracy, are shown in Table 5. The basic setting is the panoptic segmentation network we developed in earlier work (Xiang et al., 2023a), which we expanded to distinguish five semantic classes instead of only two (tree and non-tree). The other tested settings (Tables 4 and 5) are obtained by applying a single modification to the base setting.
Table 4: Δ=Δ Δ=Δ Δ=Δ Δ=Δ Instance segmentation and semantic segmentation results under different settings. In each column, bold font indicates the best results.
@@
Completeness Omission error Commission error F-score Cov oAcc mAcc mIoU
Basic setting 79.3 20.8 21.2 79.0 77.0 92.6 81.2 73.0
+ binary semantic loss 83.0 17.0 21.6 80.6 79.8 92.5 80.4 72.4
+ class weights 78.9 21.1 22.5 78.2 77.1 90.9 83.9 70.7
+ height weights 80.8 19.2 21.4 79.7 78.0 91.8 78.4 70.4
+ region weights 80.5 19.5 20.0 80.2 77.4 92.3 79.0 71.4
+ intensity 78.9 21.1 14.7 82.0 76.7 93.0 78.9 72.7
+ return number 81.7 18.3 13.5 84.1 78.2 92.9 81.6 73.7
+ scan angle rank 81.7 18.3 18.5 81.6 79.0 93.2 81.8 74.3
+ hand-crafted features 81.4 18.6 18.1 81.7 78.9 93.6 82.6 75.7
+ elastic distortion and Δ=Δ Δ=Δ Δ=Δ Δ=Δ subsampling 83.3 16.7 16.5 83.4 79.5 92.6 80.5 72.4
+ TreeMix 82.0 18.0 11.7 85.1 78.1 93.0 80.8 73.5
@@
+ hand-crafted features 89.1 78.7 55.6 94.3 60.7 75.7 94.2 97.3 95.7
+ elastic distortion and Δ=Δ Δ=Δ Δ=Δ Δ=Δ subsampling 85.8 73.9 50.8 93.9 57.8 72.4 92.6 96.6 94.6
@@
The tested setting can be categorized into three groups. The first group modifies the loss function or the sampling of training examples, i.e., the available input data is not modified, but used in a different manner. One way to do this is to add explicit supervision for the binary tree vs. non-tree task. Δ=Δ Δ=Δ Δ=Δ Δ=Δ It is implemented by adding a fully connected layer after the five-class semantic segmentation branch, and supervised with the binary classification loss. At first glance this may appear superfluous, since the binary reference is derived by simply aggregating the fine-grained labels (stems, live branches and dead branches into the tree class, ground and low vegetation into the non-tree class). Still the aggregation adds information that otherwise is unavailable to the network, namely the hierarchical structure of the label space. E.g., as soon as a point is part of a tree, it can no longer be confused with the low vegetation, hence the classifier can invest its capacity into difficult decisions within the tree class, such as separating live from dead branches. Indeed, we empirically find that adding the binary labels increases the IoU for tree vs. non-tree by 1.8 percent points (Table 5). A more correct set of tree points, in turn, better supports instance discrimination, so that the F-score for individual tree segmentation also improves by 1.6 percent points (see Table 4). The price to pay is a slightly higher confusion between the non-tree classes ground and low vegetation.
Also in this group fall different sampling and reweighting strategies. Sampling based on class frequency significantly improves the segmentation of the under-represented stem class (+3.4 pp), and also of the dead branches (+1.2 pp). However the associated accuracy loss for the ground (--2.3 pp) hurts overall performance. Reweighting based on point height is meant to compensate the sampling bias of the ALS recording geometry, vertically through the canopy. Reweighting Δ=Δ Δ=Δ Δ=Δ Δ=Δ by the regional class distribution is a coarser version of class balancing, based on regional forest composition. The latter two strategies, however, both result in a moderate decline in IoU, despite small gains in regions with little training data.
@@
The third group are tree-specific data augmentation strategies to synthetically increase the size and diversity of the training set. We have experimented with two approaches: elastic distortion with Δ=Δ Δ=Δ Δ=Δ Δ=Δ subsampling, and TreeMix. The idea behind the use of elastic distortion is to better account for curved and non-vertical stems, which are too rare to be learned well, but may still occur in the test data. The reasoning behind Δ=Δ Δ=Δ Δ=Δ Δ=Δ subsampling, i.e., removing points from the training examples, is to simulate points lost due to occlusion, scattering, etc., and make the network more robust against missing data. These low-level augmentation methods increase the instance segmentation F-score by 4.4 pp, but slightly reduce the semantic segmentation performance. TreeMix is a more informed augmentation strategy, where samples are synthesised by mixing semantically meaningful entities (i.e., trees) from different samples. It achieves the best single tree segmentation result of 85.1% F-score (+6.1 pp), while also improving semantic segmentation.
@@
Figure 4: Δ=Δ Δ=Δ Δ=Δ Δ=Δ Confusion matrix between the five semantic categories, computed over all points of the test set.
@@
In addition, NIBIO, CULS, and SCION are pure coniferous or conifer-dominated forests. Coniferous trees are generally straighter and have slimmer crowns with smaller branches. On the contrary, RMIT is native dry sclerophyll Eucalypt forest, where eucalyptus trees have few, but thick branches and expansive crowns. TUWIEN is a deciduous-dominated alluvial forest, where trees have many branches and low crowns, varying greatly in size. Expectedly, our results show that segmenting tall and straight, slender trees is easier; even when they are located close to each other (Table 6 and Figure 5). Whereas for trees with dispersed crowns and complex branch structure, individual tree segmentation is more challenging. A further reason for this discrepancy the limited diversity of training data for the forest characteristics of RMIT and TUWIEN, which each only have a single plot of training data Δ=Δ Δ=Δ Δ=Δ Δ=Δ (Figure 2).
Table 6: Δ=Δ Δ=Δ Δ=Δ Δ=Δ Individual tree segmentation results for each plot in the test set.
@@
(Commission Δ=Δ Δ=Δ Δ=Δ Δ=Δ error)
F-score Cov
1 (CULS) 20 20 23 00000020 (100.0%) 000000 (0.0%) 0000003 (13.0%) 93.0% 98.2%
2 (NIBIO) 37 26 26 26 (70.3%) 11 (29.7%) 0 (0.0%) 82.5% 62.4%
3 (NIBIO) 30 23 29 29 (96.7%) 1 (3.3%) 0 (0.0%) 98.3% 88.3%
4 (NIBIO) 27 21 26 26 (96.3%) 1 (3.7%) 0 (0.0%) 98.1% 84.6%
5 (NIBIO) 20 15 18 18 (90.0%) 2 (10.0%) 0 (0.0%) 94.7% 84.4%
6 (NIBIO) 28 18 29 26 (92.9%) 2 (7.1%) 3 (10.3%) 91.2% 81.7%
7 (NIBIO) 19 19 18 16 (84.2%) 3 (15.8%) 2 (11.1%) 86.5% 75.2%
8 (RMIT) 64 27 54 41 (64.1%) 23 (35.9%) 13 (24.1%) 69.5% 60.6%
9 (SCION) 25 25 25 23 (92.0%) 2 (8.0%) 2 (8.0%) 92.0% 86.7%
10 (SCION) 18 18 15 15 (83.3%) 3 (16.7%) 0 (0.0%) 90.9% 79.4%
11 (TUWIEN) 35 35 37 25 (71.4%) 10 (28.6%) 12 (32.4%) 69.4% 58.3%